# Language as Prior, Vision as Calibration: Metric Scale Recovery for Monocular Depth Estimation

### v1

@@ I Introduction
Monocular depth estimation, which infers scene depth from a single RGB image, is a key component in 3D perception pipelines [thorough_review, Li2023]. However, metric scale is inherently unobservable from a single view, making metric depth estimation an ill-posed problem. Although modern foundation models yield robust relative depth estimates that transfer effectively across datasets [yang2024depthanythingv2], the conversion to absolute metric depth remains contingent upon resolving a global scale ambiguity, which can be accurately modeled as an affine transformation in inverse depth space [eigen2014depth]. This ambiguity becomes especially problematic under domain shift, where even minor calibration errors can lead to significant scale drift and degrade downstream task performance.
Existing solutions for metric depth estimation typically fall into two categories. The first paradigm leverages auxiliary sensing or geometric cues (e.g., sparse range measurements, multi-view constraints, or acquisition-time metadata) to anchor absolute scale at capture or inference time [lin2025prompting], but these signals are often unavailable for unconstrained in-the-wild imagery. The second paradigm learns metric scale through supervised training, typically by training on datasets with ground-truth depth conditioned on known camera intrinsics [bhat2023zoedepth]. However, the learned scale estimates often exhibit strong dependence on the training domain and sensor configuration, leading to poor generalization under shifts in depth range, camera parameters, or scene statistics. As modern vision systems increasingly operate on diverse and heterogeneous RGB collections without consistent sensing conditions, the disconnect between widely available images and reliable metric grounding remains a critical bottleneck [zeng2024rsa].
To bridge this gap without additional sensors, recent work has explored vision–language pretraining as a source of scene-level priors for metric scale. Intuitively, language can convey coarse but informative world-scale hints: reference to object categories (e.g., car vs toy car) and scene types (e.g., indoor rooms vs. urban streets) correlate with typical object sizes and camera-to-scene distances, which constrains the plausible order of magnitude of metric depth. Large-scale vision–language models trained on image-text pairs, such as CLIP [radford2021learning], implicitly encode knowledge about object semantics with their typical sizes, and contextual scene structures that inform coarse metric scale. Prior studies leverage CLIP-style representations for depth estimation, including zero-shot depth via CLIP responses [zhang2022can] and prompting or task conditioning for improved transfer [auty2023learning]. In parallel, WorDepth [zeng2024wordepth] leverages captions as a probabilistic prior to constrain plausible 3D scene structure, while RSA [zeng2024rsa] uses language to regress global scale and shift for calibrating a frozen relative-depth backbone to metric depth under an affine-in-inverse-depth model.
@@
This motivates a role-aware design. Vision provides geometric structure, while language supplies a coarse, uncertainty-aware prior over global metric calibration. We study metric grounding under a frozen-backbone protocol: both the relative-depth backbone and the CLIP text encoder are kept fixed, and only lightweight calibration heads are trained. This setup preserves the backbone’s transferable relative geometry and isolates the generalization of global calibration under domain shift. Building on this protocol, we propose a language-guided depth calibration framework that operates atop a frozen relative-depth backbone. Specifically, we introduce two lightweight and complementary heads. A caption-conditioned head predicts an uncertainty-aware envelope that bounds feasible scale-and-shift parameters in an unconstrained calibration space. A vision-conditioned selector aggregates frozen multi-scale visual features to choose an image-specific calibration within this envelope. During training, we compute a closed-form per-image least-squares calibration in inverse depth as an oracle signal to supervise the calibration heads. We train the envelope predictor and selector jointly with metric depth supervision. Trained on NYUv2 and KITTI, our calibration improves in-domain performance and shows stronger zero-shot robustness on SUN-RGBD and DDAD than language-guided baselines [silberman2012indoor, geiger2013vision, song2015sun].
@@ III-A Experimental setup
Datasets. We use NYUv2 [silberman2012indoor] (indoor), KITTI [eigen2014depth, geiger2013vision] (outdoor), and additionally evaluate on SUN-RGBD [song2015sun] and DDAD [guizilini20203d]. NYUv2 consists of $480\times 640$ indoor images with metric depth, using the standard train/test split. KITTI consists of $352\times 1216$ outdoor driving images; we adopt the Eigen split [eigen2014depth] and follow common practice to exclude test frames without valid LiDAR ground truth. SUN-RGBD and DDAD are used for cross-dataset evaluation and provide 5,050 test images and 3,950 validation images, respectively.
Evaluation Metrics. We report standard metric depth metrics [eigen2014depth]: Abs Rel, RMSE, $\mathrm{RMSE}_{\log}$, $\log_{10}$, and threshold accuracies $\delta<1.25^{k}$.
Depth Models. We evaluate two representative depth foundation models: DPT [ranftl2021vision] and DepthAnything [yang2024depthanythingv2]. We use DPT-Hybrid fine-tuned on NYUv2 and KITTI, and DepthAnything-Small (24.8M parameters). Following our problem setting, we adopt a unified global scale-and-shift calibration to convert relative depth to metric depth, instead of using dataset-specific pixel-wise metric decoders (e.g., ZoeDepth-style [bhat2023zoedepth]). All baselines are re-implemented under this protocol, which reflects deployment scenarios with frozen backbones and limited metric supervision.
@@ III-B Implementation Details
Caption protocol. Since standard depth benchmarks do not provide paired text, we generate captions using LLaVA v1.6 [liu2024improved]. We pre-generate K=15K{=}15 one-sentence captions per image using LLaVA v1.6 with Vicuna and Mistral checkpoints, with 5 prompt templates per model. During training, we randomly sample one caption per iteration. During testing, we generate one caption per image using a fixed prompt template and decoding configuration for reproducibility. To reduce sensitivity to any single caption generator, we train with multi-caption augmentation and explicitly evaluate caption-induced variance in Table VI.
Hyperparameters. We use AdamW [loshchilov2019decoupled] and a cosine learning-rate schedule, where the learning rate decays from $3\times 10^{-5}$ to $1\times 10^{-5}$ over 50 epochs. We use a batch size of 8 and set the envelope radius upper bound to $r_{\max}=3$ in all experiments. For numerical stability, we set $\varepsilon$ to $10^{-6}$. The regularization weight $\lambda_{r}$ is selected via log-scale grid search on the validation split of the training dataset(s) within $10^{-3},10^{-1}]$, and fix it for all experiments. Unless otherwise specified, we set $\lambda_{\mathrm{env}}=0.1$ and $\lambda_{\mathrm{cal}}=1.0$ to balance the structural prior and calibration distillation. We use the same small constant $\varepsilon$ for all numerical clamps throughout the paper. All experiments are conducted on a single NVIDIA RTX 3090 GPU (24GB).

### v2

@@ I Introduction
Monocular depth estimation, which infers scene depth from a single RGB image, is a key component in 3D perception pipelines [1, 7]. However, metric scale is inherently unobservable from a single view, making metric depth estimation an ill-posed problem. Although modern foundation models yield robust relative depth estimates that transfer effectively across datasets [15], the conversion to absolute metric depth remains contingent upon resolving a global scale ambiguity, which can be accurately modeled as an affine transformation in inverse depth space [4]. This ambiguity becomes especially problematic under domain shift, where even minor calibration errors can lead to significant scale drift and degrade downstream task performance.
Existing solutions for metric depth estimation typically fall into two categories. The first paradigm leverages auxiliary sensing or geometric cues (e.g., sparse range measurements, multi-view constraints, or acquisition-time metadata) to anchor absolute scale at capture or inference time [8], but these signals are often unavailable for unconstrained in-the-wild imagery. The second paradigm learns metric scale through supervised training, typically by training on datasets with ground-truth depth conditioned on known camera intrinsics [3]. However, the learned scale estimates often exhibit strong dependence on the training domain and sensor configuration, leading to poor generalization under shifts in depth range, camera parameters, or scene statistics. As modern vision systems increasingly operate on diverse and heterogeneous RGB collections without consistent sensing conditions, the disconnect between widely available images and reliable metric grounding remains a critical bottleneck [17].
To bridge this gap without additional sensors, recent work has explored vision–language pretraining as a source of scene-level priors for metric scale. Intuitively, language can convey coarse but informative world-scale hints: reference to object categories (e.g., car vs toy car) and scene types (e.g., indoor rooms vs. urban streets) correlates with typical object sizes and camera-to-scene distances, which constrains the plausible order of magnitude of metric depth. Large-scale vision–language models trained on image-text pairs, such as CLIP [11], implicitly encode knowledge about object semantics with their typical sizes, and contextual scene structures that inform coarse metric scale. Prior studies leverage CLIP-style representations for depth estimation, including zero-shot depth via CLIP responses [18] and prompting or task conditioning for improved transfer [2]. In parallel, WorDepth [16] leverages captions as a probabilistic prior to constrain plausible 3D scene structure, while RSA [17] uses language to regress global scale and shift for calibrating a frozen relative-depth backbone to metric depth under an affine-in-inverse-depth model.
@@
This motivates a role-aware design. Vision provides geometric structure, while language supplies a coarse, uncertainty-aware prior over global metric calibration. We study metric grounding under a frozen-backbone protocol: both the relative-depth backbone and the CLIP text encoder are kept fixed, and only lightweight calibration heads are trained. This setup preserves the backbone’s transferable relative geometry and isolates the generalization of global calibration under domain shift. Building on this protocol, we propose a language-guided depth calibration framework that operates atop a frozen relative-depth backbone. Specifically, we introduce two lightweight and complementary heads. A caption-conditioned head predicts an uncertainty-aware envelope that bounds feasible scale-and-shift parameters in an unconstrained calibration space. A vision-conditioned selector aggregates frozen multi-scale visual features to choose an image-specific calibration within this envelope. During training, we compute a closed-form per-image least-squares calibration in inverse depth as an oracle signal to supervise the calibration heads. We train the envelope predictor and selector jointly with metric depth supervision. Trained on NYUv2 and KITTI, our calibration improves in-domain performance and shows stronger zero-shot robustness on SUN-RGBD and DDAD than language-guided baselines [13, 5, 14].
@@ III-A Experimental setup
Datasets. We use NYUv2 [13] (indoor), KITTI [4, 5] (outdoor), and additionally evaluate on SUN-RGBD [14] and DDAD [6]. NYUv2 consists of $480\times 640$ indoor images with metric depth, using the standard train/test split. KITTI consists of $352\times 1216$ outdoor driving images; we adopt the Eigen split [4] and follow common practice to exclude test frames without valid LiDAR ground truth. SUN-RGBD and DDAD are used for cross-dataset evaluation and provide 5,050 test images and 3,950 validation images, respectively.
Evaluation Metrics. We report standard metric depth metrics [4]: Abs Rel, RMSE, $\mathrm{RMSE}_{\log}$, $\log_{10}$, and threshold accuracies $\delta<1.25^{k}$.
Depth Models. We evaluate two representative depth foundation models: DPT [12] and DepthAnything [15]. We use DPT-Hybrid fine-tuned on NYUv2 and KITTI, and DepthAnything-Small (24.8M parameters). Following our problem setting, we adopt a unified global scale-and-shift calibration to convert relative depth to metric depth, instead of using dataset-specific pixel-wise metric decoders (e.g., ZoeDepth-style [3]). All baselines are re-implemented under this protocol, which reflects deployment scenarios with frozen backbones and limited metric supervision.
@@ III-B Implementation Details
Caption protocol. Since standard depth benchmarks do not provide paired text, we generate captions using LLaVA v1.6 [9]. We pre-generate K=15K{=}15 one-sentence captions per image using LLaVA v1.6 with Vicuna and Mistral checkpoints, with 5 prompt templates per model. During training, we randomly sample one caption per iteration. During testing, we generate one caption per image using a fixed prompt template and decoding configuration for reproducibility. To reduce sensitivity to any single caption generator, we train with multi-caption augmentation and explicitly evaluate caption-induced variance in Table VI.
Hyperparameters. We use AdamW [10] and a cosine learning-rate schedule, where the learning rate decays from $3\times 10^{-5}$ to $1\times 10^{-5}$ over 50 epochs. We use a batch size of 8 and set the envelope radius upper bound to $r_{\max}=3$ in all experiments. For numerical stability, we set $\varepsilon$ to $10^{-6}$. The regularization weight $\lambda_{r}$ is selected via log-scale grid search on the validation split of the training dataset(s) within $10^{-3},10^{-1}]$, and fix it for all experiments. Unless otherwise specified, we set $\lambda_{\mathrm{env}}=0.1$ and $\lambda_{\mathrm{cal}}=1.0$ to balance the structural prior and calibration distillation. We use the same small constant $\varepsilon$ for all numerical clamps throughout the paper. All experiments are conducted on a single NVIDIA RTX 3090 GPU (24GB).